OpenStudy (anonymous):

I need help creating a quadratic function with two real zeros. Will give medal to best answer (:

OpenStudy (anonymous):

Oops + fan

OpenStudy (anonymous):

Choose your two favourite real numbers. Think of them as 'a' and 'b'. Then your quadratic function is (x-a)(x-b).

OpenStudy (anonymous):

I thought a function was F(x)

OpenStudy (jdoe0001):

as BillBell said suggested.... do like so

OpenStudy (anonymous):

I am confused though :/

OpenStudy (jdoe0001):

ok.... have you covered quadratics yet?

OpenStudy (anonymous):

Yeah the unit I am on is actually called Quadratic functions but my lessons aren't exactly well on explaining

OpenStudy (jdoe0001):

ok... well. lemme .. pick a quick example \(\large {y=x^2+3x-4\implies 0=x^2+3x-4\implies 0=(x+4)(x-1) \\ \quad \\ meaning\implies (x+4)(x-1)\implies \begin{cases} x+4=0\implies x=-4 \\ \quad \\ x-1=0\implies x=1 \end{cases} }\) so if for example, you pick 2 values for "x", any values really, as BillBell suggested, let's say "a" and "b", which could be anyting then \(\large \begin{cases} x=a\implies x-a=0 \\ \quad \\ x=b\implies x-b=0 \end{cases}\implies \begin{array}{llll} (x-a)(x-b)=0 \\ \quad \\ (x-a)(x-b)=\textit{original function} \end{array} \)

OpenStudy (jdoe0001):

and "a" and "b" can be pretty much anything, then just do a FOIL and that'd give the origiinal function

OpenStudy (anonymous):

So If i choose (x-2)(x-4) My answer would be x^2-4-2x-8? I am sorry if its wrong or if i got the signs wrong

OpenStudy (jdoe0001):

\(\bf (x-2)(x-4)\implies x^2-4x-2x+8\implies x^2-2x+8\) yes

OpenStudy (anonymous):

Ohh i see I forgot to put the x after the 4 and how did you get the +8?

OpenStudy (anonymous):

To answer your question about what a function is, there are a few ways of thinking about them. One simple way is that a function takes numbers and converts them into other numbers. For example, (x-2)(x-4) is a function in that it can take the number 2 and convert it to 0; it takes the number 3 and converts it to 1. In fact, this is a function that can convert any number (that you are likely to think of) into another number. If you wanted to give it a name for convenience you _could_ call it F(x) but this name wouldn't really tell you much about the function.

OpenStudy (anonymous):

\[( x - 2 ) ( x - 4 )\] expanding the first part \[= x ( x - 4 ) - 2 ( x - 4 )\] multiplying out the first parenthesised expression \[= x ^{2} - 4x - 2 ( x - 4 )\] now multiplying out the second one \[= x ^{2} - 4x - 2x + 8\] collect all the x-squared terms (just one), all the x terms (two), all the number terms (one) \[= x ^{2} - 6x + 8\] And, mercifully, we're finished.